# Fidelity and criticality in the nonreciprocal Aubry-André-Harper model

### v2

@@ V Conclusion
In the ground state of the AAH model, characterized by real eigenenergies, we observe that the scaling behavior of the fidelity susceptibilities is perfectly consistent with that of the Hermitian AAH model, exhibiting an equivalent correlation length critical exponent of $\nu=1$. In the first-excited state, where the system has a $\mathcal{PT}$ transition, the fidelity susceptibilities exhibit different scaling laws for the odd/even sites. For even lattices, the self-normalized fidelity susceptibilities exhibit a scaling behavior of $N^{2}$ akin to that observed in the ground state fidelity susceptibilities. In contrast, on odd lattices, the self-normalized fidelity susceptibilities follow a linear scaling law, $\chi_{F_{1}}^{S}(\lambda_{m})\propto N$, while the biorthogonal fidelity susceptibilities diverge. Exploring whether these scaling laws persist in the many-body AAH model would be an intriguing avenue for future investigation.
@@
Note added.- After completing our work, we became aware of a related research [79] focusing on characterizing the phase transition of a generalized AAH model using fidelity susceptibility.
Acknowledgements.
@@
G.S. is appreciative of support from "the Fundamental Research Funds for the Central Universities, under the Grant No. NS2023055", the NSFC under the Grants No. 11704186, and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics. C.Z., Z.C. and G.W. are appreciative of support from the Innovation Training Fund for College Students under Grant No. 202310287211Y.

### v3

@@ Abstract
In the ground state of the AAH model, characterized by real eigenenergies, we observe that the scaling behavior of the fidelity susceptibilities is perfectly consistent with that of the Hermitian AAH model, exhibiting an equivalent correlation length critical exponent of $\nu=1$. In the first-excited state, the fidelity susceptibilities exhibit different scaling laws for the odd/even sites. For even lattices, the self-normalized fidelity susceptibilities exhibit a scaling behavior of $N^{2}$ akin to that observed in the ground state fidelity susceptibilities. In contrast, on odd lattices, where the system has a $\mathcal{PT}$ transition, the self-normalized fidelity susceptibilities follow a linear scaling law, $\chi_{F_{1}}^{S}(\lambda_{m})\propto N$, while the biorthogonal fidelity susceptibilities diverge. The $1/2$ value of the biorthogonal fidelity has been observed in other systems Tu et al. (2023). Exploring whether these scaling laws persist in the many-body AAH model and other systems would be an intriguing avenue for future investigation.
@@
Note added.- After completing our work, we became aware of two related works focusing on characterizing the phase transition of a generalized AAH model using fidelity susceptibility Ren et al. (2024) and the scaling theory of the non-Hermitian disorder AAH model Sun et al. (2024), respectively.
@@
Acknowledgments.- G.S. is appreciative of support from "the Fundamental Research Funds for the Central Universities, under the Grant No. NS2023055", the NSFC under the Grants No. 11704186, and the High Performance Computing Platform of Nanjing University of Aeronautics and Astronautics. C.Z., Z.C. and G.W. are appreciative of support from the Innovation Training Fund for College Students under Grant No. 202310287211Y.